# GenPO: Generative Diffusion Models Meet On-Policy Reinforcement Learning

### v2

@@ Abstract
Recent advances in reinforcement learning (RL) have demonstrated the powerful exploration capabilities and multimodality of generative diffusion-based policies. While substantial progress has been made in offline RL and off-policy RL settings, integrating diffusion policies into on-policy frameworks like PPO remains underexplored. This gap is particularly significant given the widespread use of large-scale parallel GPU-accelerated simulators, such as IsaacLab, which are optimized for on-policy RL algorithms and enable rapid training of complex robotic tasks. A key challenge lies in computing state-action log-likelihoods under diffusion policies, which is straightforward for Gaussian policies but intractable for flow-based models due to irreversible forward-reverse processes and discretization errors (e.g., Euler-Maruyama approximations). To bridge this gap, we propose GenPO, a generative policy optimization framework that leverages exact diffusion inversion to construct invertible action mappings. GenPO introduces a novel doubled dummy action mechanism that enables invertibility via alternating updates, resolving log-likelihood computation barriers. Furthermore, we also use the action log-likelihood for unbiased entropy and KL divergence estimation, enabling KL-adaptive learning rates and entropy regularization in on-policy updates. Extensive experiments on eight IsaacLab benchmarks, including legged locomotion (Ant, Humanoid, Anymal-D, Unitree H1, Go2), dexterous manipulation (Shadow Hand), aerial control (Quadcopter), and robotic arm tasks (Franka), demonstrate GenPO’s superiority over existing RL baselines. Notably, GenPO is the first method to successfully integrate diffusion policies into on-policy RL, unlocking their potential for large-scale parallelized training and real-world robotic deployment.
@@ 1 Introduction
In recent years, generative diffusion models, such as DDPM ho2020denoising , DDIM song2020denoising , and flow matching liu2022flow ; lipmanflow , have attracted considerable attention from reinforcement learning researchers due to their powerful exploration capabilities and multimodality compared with traditional parameterized Gaussian policies. However, existing works in this field almost merely focus on offline RL wang2022diffusion ; Ada_2024 ; kang2024efficient ; chen2022offline ; chen2023boosting ; hansen2023idql ; zhu2023madiff ; ni2023metadiffuser and off-policy RL yang2023policy ; ding2024diffusion ; psenka2023learning ; park2025flow ; wang2024diffusion methods, and the integration of diffusion or flow-based policies into on-policy reinforcement learning like PPO schulman2017proximal remains largely unexplored. However, existing massively parallel GPU-accelerated simulators, such as Isaac Gym makoviychuk2021isaac and its successor, IsaacLab, mainly li2023parallel benefit the training of on-policy reinforcement learning algorithms, especially PPO, which are capable of achieving substantial performance in a relatively short wall-clock time. This has resulted in a fundamental dilemma, hindering the practical deployment of existing diffusion/flow-based RL algorithms in real-world applications chi2023diffusion ; li2023crossway ; pearce2023imitating ; reuss2023goal ; ding2025diopt .
@@
Different from diffusion models ding2024diffusion applied in off-policy RL, the primary challenge in integrating flow policies with on-policy reinforcement learning is how to compute the log-likelihood of state-action pairs, while it is straightforward for Gaussian policies. The underlying reason here lies in the inconsistency and non-reversibility between the forward and reverse processes of the generative diffusion policy, caused by the Euler-Maruyama (EM) discretization sarkka2019applied ; celik2025dime . Nevertheless, the optimization objective in on-policy RL necessitates the probability density of the given state-action pair. In this context, estimating the probability density of a given state-action pair in flow or diffusion models is nontrivial, and further utilizing this estimate for diffusion or flow policy improvement in the on-policy RL paradigm presents additional challenges.
@@
To address these challenges, we propose GenPO, motivated by the exact diffusion inversion wallace2023edict , which constructs an invertible flow mapping and thus avoids the mismatch of forward and reverse processes of the diffusion policy. Specifically, we first construct a new Markov process with the action space twice the size of the original in GenPO due to the limitation of exact diffusion inversion, and then enable reversible flow processes by alternately updating the two parts of the dummy action. In that case, we can calculate the exact probability density of the given action in this new Markov decision problem via change of variables theory like normalizing flow rezende2015variational ; dinh2016density . Moreover, the two parts of the dummy action will be averaged as one and then mapped to the original action space during inference. Hence, we can perform GenPO to update the generative diffusion policy in the reformulated Markov decision problem, but finally obtain the optimal solution to the original task.
@@
In addition, with the exact probability density, we estimate cover1999elements the entropy of the diffusion policy and the Kullback-Leibler (KL) divergence with respect to the behavior diffusion policy. This allows for the combination of entropy regularization and KL-adaptive learning rate adjustment into GenPO, which has been confirmed effective in PPO. To demonstrate the superiority of GenPO, we conduct experiments on 8 IsaacLab benchmarks makoviychuk2021isaac , which cover robot ant, humanoid, quadcopter, Franka robot arm haddadin2024franka , Shadow dexterous hand tuffield2003shadow , ANYbotics anymal-D, and Unitree legged robots zhuang2024humanoid . The final results show that the proposed GenPO outperforms previous RL baselines in cumulative returns, while maintaining comparable sample efficiency and faster convergence. It is worth noting that GenPO is the first trial to incorporate the generative diffusion policy into on-policy RL algorithms and opens brand-new avenues for diffusion-type policies applied in large-scale parallel simulators and real-world robot control. Our contributions are summarized as follows:
@@ 2 Related Works
In this section, we review the literature of generative models such as VAE kingma2013auto , GAN creswell2018generative , normalizing flow rezende2015variational , diffusion zhang2021diffusion , and flow-based model liu2022flow ; lipman2022flow for policy learning, and generally divide them into two classes according to their learning paradigm.
@@
Generative Policy for Online Reinforcement Learning. The goal of online RL sutton2018reinforcement is to learn an optimal policy with the interactions of the given environment. It is challenging to train a generative model policy within an online RL paradigm due to the absence of the action label. Generative policy learning methods in online RL can be broadly grouped into two paradigms. The first paradigm views the generative policies as a black-box function approximation like neural networks, and optimizes them via the policy gradient. In this paradigm, representative methods include DACER wang2024diffusion , CPQL chen2024boosting , FlowPG brahmanage2023flowpg , SAC-NF mazoure2020leveraging , and TRPO-NF tang2018boosting , which directly apply deterministic or policy gradient on the generative model policies. By contrast, the second category leverages the internal structure and mechanisms of the generative model itself to perform the policy improvement. For instance, DIPO yang2023policy and QVPO ding2024diffusion respectively utilize the gradient and the value magnitude of the Q-function to identify optimal actions, which are then applied to the variational loss of the diffusion model. Additionally, QSM psenka2023learning and MaxEntDP dong2025maximum regard the noise network in the diffusion model as the score function of the target policy distribution and then employ the gradient of the Q-function and its variants to train the noise network. As to normalizing flow–based approaches, MEow chao2024maximum exploits the layer-wise nonlinear property of the flow network to train the Q and V functions, which yields the corresponding maximum-entropy normalizing flow policy based on SAC haarnoja2018soft ; haarnoja2018soft2 . Overall, current approaches that employ generative-model policies in online RL have predominantly focused on integrating these models with off-policy RL algorithms like SAC, with few efforts to combine them with on-policy RL methods.
@@
Generative Policy for Other Learning Paradigms. Compared with generative policy for online RL, more existing research works on generative policy have been devoted to offline RL levine2020offline and imitation learning paradigms, wherein policies are learned from offline datasets without environment interaction. For offline RL, BCQ fujimoto2019off , PLAS zhou2021plas , SPOT wu2022supported adopt VAE kingma2013auto policy regularizer to prevent the policy from optimizing outside the support of the dataset. Besides, CPED zhang2023constrained utilizes flow GAN grover2017flow to achieve more accurate policy regularization. Furthermore, diffusion-QL wang2022diffusion , EDP kang2024efficient , SRDP Ada_2024 , CEP lu2023contrastive , and FQL park2025flow leverage diffusion ho2020denoising and flow liu2022flow ; lipman2022flow policies to accurately model the policy distribution in offline datasets, thereby obtaining policies with superior generalization capability. Regarding imitation learning, diffusion policy chi2023diffusion and diffuser janner2022planning serve as canonical examples of using diffusion models to fit the state–action mapping and state–trajectory mapping, respectively. The latter one is also referred to as the diffusion planner. Subsequently, FMIL rouxel2024flow , AVDC ko2023learning , and decision diffuser ajay2022conditional follow their steps and demonstrate enhanced performance in different robots. Furthermore, DPPO ren2024diffusion applies policy gradient to fine-tune the diffusion policy trained with the offline dataset. Nonetheless, DPPO merely explores the use of the RL technique in finetuning an offline-pretrained diffusion policy, rather than representing a diffusion-based online RL algorithm.
@@ 3.1 On-policy Reinforcement Learning
Reinforcement learning problems sutton2018reinforcement are formalized as Markov decision processes (MDPs), defined by the tuple $(\mathcal{S},\mathcal{A},p,r,\rho_{0},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $p(s^{\prime}\mid s,a):\mathcal{S}\times\mathcal{S}\times\mathcal{A}\rightarrow[0,\infty)$ denotes the transition dynamics, $r(s,a):\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\rho_{0}(s):\mathcal{S}\rightarrow[0,\infty)$ is the distribution of the initial state, and $\gamma\in[0,1)$ is the discount factor for the value estimation. The goal of RL is to maximize the expected discounted return as $J(\theta)\;=\;\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$, where $\pi_{\theta}(a\!\mid\!s)$ is the behavior of the agent parametrized by $\theta$, and induces a corresponding distribution over behavior trajectories $\tau$. For convenience, RL also defines two different value functions: the state value function $V_{\pi}({s})=\mathbb{E}_{\tau\sim\pi}[\sum_{t=0}\gamma^{t}r_{t}|{s}_{0}={s}]$ and the state-action value function $Q_{\pi}(s,a)=\mathbb{E}_{\tau\sim\pi}[\sum_{t=0}\gamma^{t}r_{t}|{a}_{0}={a},{s}_{0}={s}]$. Besides, the advantage function, which represents the benefit of taking one specific action compared with the expected return at the given state, is defined as $A_{\pi}({s},{a})=Q_{\pi}({s},{a})-V_{\pi}({s})$.
On-policy RL algorithms train the policy with the transitions generated by the current policy itself, while off-policy RL algorithms use the transitions from a different policy as training samples. While on-policy RL algorithms preserve the consistency between the behavior and target policies, but often suffer from high variance and sample inefficiency. Proximal Policy Optimization (PPO) schulman2017proximal addresses these challenges by introducing the clipped surrogate objective:
@@
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\mathrm{old}}}(a_{t}\mid s_{t})}$ and $\hat{A}_{t}$ is the estimation of the advantage function via GAE schulman2015high . By leveraging the importance sampling technique and trust-region insights schulman2015trust without second-order complexity for approximation of the KL divergence constraint, PPO can achieve stable and efficient policy updates. The simplicity, robustness, and strong empirical performance on continuous-control benchmarks makoviychuk2021isaac ; zhuang2024humanoid of PPO have made it the first algorithm to be adopted in many domains.
@@ 3.2 Generative Diffusion Models and Flow Matching
Generative diffusion models ho2020denoising ; song2020score ; song2020denoising are a kind of powerful latent variable generative model that can transform samples from the standard Gaussian distribution $x_{T}\sim\mathcal{N}(0,I)$ to the data distribution $x_{0}\sim p_{data}$ via the denoising procedure with the noise network. The relationship between $x_{0}$ and $x_{T}$ is defined by the forward SDE of diffusion
@@
where $W_{t}$ the standard Wiener process (a.k.a., Brownian motion), $f(x_{t},t)$ is a vector-valued function called the drift coefficient of $x_{t}$, and $g(t)$is a scalar function known as the diffusion coefficient of $x_{t}$. According to song2020score , the reverse-time SDE of diffusion (i.e., denoising procedure) is
@@
where $\overline{W}_{t}$ is a standard Wiener process when time flows backwards from $T$ to 00. In practice, DDPM ho2020denoising discretizes the SDE and approximates the forward step as $q(x_{t}\mid x_{t-1}):=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$, where $\beta_{t}$ is the variance schedule. The noise network $\boldsymbol{\epsilon}_{\theta}(x_{t},t)$ in DDPM can be trained by the variational lower bound loss:
$\mathbb{E}_{t\sim[1,T],x_{0},\boldsymbol{\epsilon}_{t}}\left[||\boldsymbol{\epsilon}_{t}-\boldsymbol{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t},t)||^{2}\right],$ (4)
@@
[…]

### v3

@@ Abstract
Recent advances in reinforcement learning (RL) have demonstrated the powerful exploration capabilities and multimodality of generative diffusion-based policies. While substantial progress has been made in offline RL and off-policy RL settings, integrating diffusion policies into on-policy frameworks like PPO remains underexplored. This gap is particularly significant given the widespread use of large-scale parallel GPU-accelerated simulators, such as IsaacLab, which are optimized for on-policy RL algorithms and enable rapid training of complex robotic tasks. A key challenge lies in computing state-action log-likelihoods under diffusion policies, which is straightforward for Gaussian policies but intractable for flow-based models due to irreversible forward-reverse processes and discretization errors (e.g., Euler-Maruyama approximations). To bridge this gap, we propose GenPO, a generative policy optimization framework that leverages exact diffusion inversion to construct invertible action mappings. GenPO introduces a novel doubled dummy action mechanism that enables invertibility via alternating updates, resolving log-likelihood computation barriers. Furthermore, we also use the action log-likelihood for unbiased entropy and KL divergence estimation, enabling KL-adaptive learning rates and entropy regularization in on-policy updates. Extensive experiments on eight IsaacLab benchmarks, including legged locomotion (Ant, Humanoid, Anymal-D, Unitree H1, Go2), dexterous manipulation (Shadow Hand), aerial control (Quadcopter), and robotic arm tasks (Franka), demonstrate GenPO’s superiority over existing RL baselines. Notably, GenPO is the first method to successfully integrate diffusion policies into on-policy RL, unlocking their potential for large-scale parallelized training and real-world robotic deployment. The official implementation of DiOpt is provided in https://github.com/wadx2019/genpo/.
@@ 1 Introduction
In recent years, generative diffusion models, such as DDPM Ho et al. (2020), DDIM Song et al. (2020a), and flow matching Liu et al. (2022); Lipman et al. (2023), have attracted considerable attention from reinforcement learning researchers due to their powerful exploration capabilities and multimodality compared with traditional parameterized Gaussian policies. However, existing works in this field almost merely focus on offline RL Wang et al. (2022); Ada et al. (2024); Kang et al. (2024); Chen et al. (2022, 2023); Hansen-Estruch et al. (2023); Zhu et al. (2023); Ni et al. (2023) and off-policy RL Yang et al. (2023); Ding et al. (2024); Psenka et al. (2023); Park et al. (2025); Wang et al. (2024b) methods, and the integration of diffusion or flow-based policies into on-policy reinforcement learning like PPO Schulman et al. (2017) remains largely unexplored. However, existing massively parallel GPU-accelerated simulators, such as Isaac Gym Makoviychuk et al. (2021) and its successor, IsaacLab, mainly Li et al. (2023b) benefit the training of on-policy reinforcement learning algorithms, especially PPO, which are capable of achieving substantial performance in a relatively short wall-clock time. This has resulted in a fundamental dilemma, hindering the practical deployment of existing diffusion/flow-based RL algorithms in real-world applications Chi et al. (2023); Li et al. (2023a); Pearce et al. (2023); Reuss et al. (2023); Ding et al. (2025).
@@
Different from diffusion models Ding et al. (2024) applied in off-policy RL, the primary challenge in integrating flow policies with on-policy reinforcement learning is how to compute the log-likelihood of state-action pairs, while it is straightforward for Gaussian policies. The underlying reason here lies in the inconsistency and non-reversibility between the forward and reverse processes of the generative diffusion policy, caused by the Euler-Maruyama (EM) discretization Särkkä and Solin (2019); Celik et al. (2025). Nevertheless, the optimization objective in on-policy RL necessitates the probability density of the given state-action pair. In this context, estimating the probability density of a given state-action pair in flow or diffusion models is nontrivial, and further utilizing this estimate for diffusion or flow policy improvement in the on-policy RL paradigm presents additional challenges.
@@
To address these challenges, we propose GenPO, motivated by the exact diffusion inversion Wallace et al. (2023), which constructs an invertible flow mapping and thus avoids the mismatch of forward and reverse processes of the diffusion policy. Specifically, we first construct a new Markov process with the action space twice the size of the original in GenPO due to the limitation of exact diffusion inversion, and then enable reversible flow processes by alternately updating the two parts of the dummy action. In that case, we can calculate the exact probability density of the given action in this new Markov decision problem via change of variables theory like normalizing flow Rezende and Mohamed (2015); Dinh et al. (2016). Moreover, the two parts of the dummy action will be averaged as one and then mapped to the original action space during inference. Hence, we can perform GenPO to update the generative diffusion policy in the reformulated Markov decision problem, but finally obtain the optimal solution to the original task.
@@
In addition, with the exact probability density, we estimate Cover (1999) the entropy of the diffusion policy and the Kullback-Leibler (KL) divergence with respect to the behavior diffusion policy. This allows for the combination of entropy regularization and KL-adaptive learning rate adjustment into GenPO, which has been confirmed effective in PPO. To demonstrate the superiority of GenPO, we conduct experiments on 8 IsaacLab benchmarks Makoviychuk et al. (2021), which cover robot ant, humanoid, quadcopter, Franka robot arm Haddadin (2024), Shadow dexterous hand Tuffield and Elias (2003), ANYbotics anymal-D, and Unitree legged robots Zhuang et al. (2024). The final results show that the proposed GenPO outperforms previous RL baselines in cumulative returns, while maintaining comparable sample efficiency and faster convergence. It is worth noting that GenPO is the first trial to incorporate the generative diffusion policy into on-policy RL algorithms and opens brand-new avenues for diffusion-type policies applied in large-scale parallel simulators and real-world robot control. Our contributions are summarized as follows:
@@ 2 Related Works
In this section, we review the literature of generative models such as VAE Kingma and Welling (2013), GAN Creswell et al. (2018), normalizing flow Rezende and Mohamed (2015), diffusion Zhang and Chen (2021), and flow-based model Liu et al. (2022); Lipman et al. (2022) for policy learning, and generally divide them into two classes according to their learning paradigm.
@@
Generative Policy for Online Reinforcement Learning. The goal of online RL Sutton and Barto (2018) is to learn an optimal policy with the interactions of the given environment. It is challenging to train a generative model policy within an online RL paradigm due to the absence of the action label. Generative policy learning methods in online RL can be broadly grouped into two paradigms. The first paradigm views the generative policies as a black-box function approximation like neural networks, and optimizes them via the policy gradient. In this paradigm, representative methods include DACER Wang et al. (2024b), CPQL Chen et al. (2024), FlowPG Brahmanage et al. (2023), SAC-NF Mazoure et al. (2020), and TRPO-NF Tang and Agrawal (2018), which directly apply deterministic or policy gradient on the generative model policies. By contrast, the second category leverages the internal structure and mechanisms of the generative model itself to perform the policy improvement. For instance, DIPO Yang et al. (2023) and QVPO Ding et al. (2024) respectively utilize the gradient and the value magnitude of the Q-function to identify optimal actions, which are then applied to the variational loss of the diffusion model. Additionally, QSM Psenka et al. (2023) and MaxEntDP Dong et al. (2025) regard the noise network in the diffusion model as the score function of the target policy distribution and then employ the gradient of the Q-function and its variants to train the noise network. As to normalizing flow–based approaches, MEow Chao et al. (2024) exploits the layer-wise nonlinear property of the flow network to train the Q and V functions, which yields the corresponding maximum-entropy normalizing flow policy based on SAC Haarnoja et al. (2018a, b). Overall, current approaches that employ generative-model policies in online RL have predominantly focused on integrating these models with off-policy RL algorithms like SAC, with few efforts to combine them with on-policy RL methods.
@@
Generative Policy for Other Learning Paradigms. Compared with generative policy for online RL, more existing research works on generative policy have been devoted to offline RL Levine et al. (2020) and imitation learning paradigms, wherein policies are learned from offline datasets without environment interaction. For offline RL, BCQ Fujimoto et al. (2019), PLAS Zhou et al. (2021), SPOT Wu et al. (2022) adopt VAE Kingma and Welling (2013) policy regularizer to prevent the policy from optimizing outside the support of the dataset. Besides, CPED Zhang et al. (2023) utilizes flow GAN Grover et al. (2017) to achieve more accurate policy regularization. Furthermore, diffusion-QL Wang et al. (2022), EDP Kang et al. (2024), SRDP Ada et al. (2024), CEP Lu et al. (2023), and FQL Park et al. (2025) leverage diffusion Ho et al. (2020) and flow Liu et al. (2022); Lipman et al. (2022) policies to accurately model the policy distribution in offline datasets, thereby obtaining policies with superior generalization capability. Regarding imitation learning, diffusion policy Chi et al. (2023) and diffuser Janner et al. (2022) serve as canonical examples of using diffusion models to fit the state–action mapping and state–trajectory mapping, respectively. The latter one is also referred to as the diffusion planner. Subsequently, FMIL Rouxel et al. (2024), AVDC Ko et al. (2023), and decision diffuser Ajay et al. (2022) follow their steps and demonstrate enhanced performance in different robots. Furthermore, DPPO Ren et al. (2024) applies policy gradient to fine-tune the diffusion policy trained with the offline dataset. Nonetheless, DPPO merely explores the use of the RL technique in finetuning an offline-pretrained diffusion policy, rather than representing a diffusion-based online RL algorithm.
@@ 3.1 On-policy Reinforcement Learning
Reinforcement learning problems Sutton and Barto (2018) are formalized as Markov decision processes (MDPs), defined by the tuple $(\mathcal{S},\mathcal{A},p,r,\rho_{0},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $p(s^{\prime}\mid s,a):\mathcal{S}\times\mathcal{S}\times\mathcal{A}\rightarrow[0,\infty)$ denotes the transition dynamics, $r(s,a):\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\rho_{0}(s):\mathcal{S}\rightarrow[0,\infty)$ is the distribution of the initial state, and $\gamma\in[0,1)$ is the discount factor for the value estimation. The goal of RL is to maximize the expected discounted return as $J(\theta)\;=\;\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$, where $\pi_{\theta}(a\!\mid\!s)$ is the behavior of the agent parametrized by $\theta$, and induces a corresponding distribution over behavior trajectories $\tau$. For convenience, RL also defines two different value functions: the state value function $V_{\pi}({s})=\mathbb{E}_{\tau\sim\pi}[\sum_{t=0}\gamma^{t}r_{t}|{s}_{0}={s}]$ and the state-action value function $Q_{\pi}(s,a)=\mathbb{E}_{\tau\sim\pi}[\sum_{t=0}\gamma^{t}r_{t}|{a}_{0}={a},{s}_{0}={s}]$. Besides, the advantage function, which represents the benefit of taking one specific action compared with the expected return at the given state, is defined as $A_{\pi}({s},{a})=Q_{\pi}({s},{a})-V_{\pi}({s})$.
On-policy RL algorithms train the policy with the transitions generated by the current policy itself, while off-policy RL algorithms use the transitions from a different policy as training samples. While on-policy RL algorithms preserve the consistency between the behavior and target policies, but often suffer from high variance and sample inefficiency. Proximal Policy Optimization (PPO) Schulman et al. (2017) addresses these challenges by introducing the clipped surrogate objective:
@@
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\mathrm{old}}}(a_{t}\mid s_{t})}$ and $\hat{A}_{t}$ is the estimation of the advantage function via GAE Schulman et al. (2015b). By leveraging the importance sampling technique and trust-region insights Schulman et al. (2015a) without second-order complexity for approximation of the KL divergence constraint, PPO can achieve stable and efficient policy updates. The simplicity, robustness, and strong empirical performance on continuous-control benchmarks Makoviychuk et al. (2021); Zhuang et al. (2024) of PPO have made it the first algorithm to be adopted in many domains.
@@ 3.2 Generative Diffusion Models and Flow Matching
Generative diffusion models Ho et al. (2020); Song et al. (2020b, a) are a kind of powerful latent variable generative model that can transform samples from the standard Gaussian distribution $x_{T}\sim\mathcal{N}(0,I)$ to the data distribution $x_{0}\sim p_{data}$ via the denoising procedure with the noise network. The relationship between $x_{0}$ and $x_{T}$ is defined by the forward SDE of diffusion
@@
where $W_{t}$ the standard Wiener process (a.k.a., Brownian motion), $f(x_{t},t)$ is a vector-valued function called the drift coefficient of $x_{t}$, and g(t)g(t)is a scalar function known as the diffusion coefficient of $x_{t}$. According to Song et al. (2020b), the reverse-time SDE of diffusion (i.e., denoising procedure) is
@@
where $\overline{W}_{t}$ is a standard Wiener process when time flows backwards from $T$ to 0. In practice, DDPM Ho et al. (2020) discretizes the SDE and approximates the forward step as $q(x_{t}\mid x_{t-1}):=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$, where $\beta_{t}$ is the variance schedule. The noise network $\boldsymbol{\epsilon}_{\theta}(x_{t},t)$ in DDPM can be trained by the variational lower bound loss:
[…]
@@
The flow matching model Liu et al. (2022); Lipman et al. (2022) is developed for faster generation compared with classical diffusion models like DDPM, and can be viewed as a special type of generative diffusion model Gao et al. (2024). Similar to DDPM, flow matching model learns a time-dependent vector field $v_{\theta}(x,t)$, i.e., noise network in DDPM, that transports samples along a predefined family of conditional distributions $\{p_{t}\}_{t\in[0,1]}$ satisfying $x_{1}\sim p_{\mathrm{data}}$ and $x_{0}\sim\mathcal{N}(0,I)$. Hence, the sample from the target distribution $p_{data}$ can be achieved via solving the ode $x_{1}=x_{0}+\int_{0}^{1}v(x_{t},t)dt$.